# GraphAllocBench: A Flexible Benchmark for Preference-Conditioned Multi-Objective Policy Learning

### v1

@@ Abstract
Preference-Conditioned Policy Learning (PCPL) in Multi-Objective Reinforcement Learning (MORL) aims to approximate diverse Pareto-optimal solutions by conditioning policies on user-specified preferences over objectives. This enables a single model to flexibly adapt to arbitrary trade-offs at run-time by producing a policy on or near the Pareto front. However, existing benchmarks for PCPL are largely restricted to toy tasks and fixed environments, limiting their realism and scalability. To address this gap, we introduce GraphAllocBench, a flexible benchmark built on a novel graph-based resource allocation sandbox environment inspired by city management, which we call CityPlannerEnv. GraphAllocBench provides a rich suite of problems with diverse objective functions, varying preference conditions, and high-dimensional scalability. We also propose two new evaluation metrics – Proportion of Non-Dominated Solutions (PNDS) and Ordering Score (OS) – that directly capture preference consistency while complementing the widely used hypervolume metric. Through experiments with Multi-Layer Perceptrons (MLPs) and graph-aware models, we show that GraphAllocBench exposes the limitations of existing MORL approaches and paves the way for using graph-based methods such as Graph Neural Networks in complex, high-dimensional combinatorial allocation tasks. Beyond its predefined problem set, GraphAllocBench enables users to flexibly vary objectives, preferences, and allocation rules, establishing it as a versatile and extensible benchmark for advancing PCPL. Code: https://github.com/jzh001/GraphAllocBench
@@ 2.1 Pareto Front
A solution $x^{*}\in X$ is called Pareto optimal if there is no other solution $x\in X$ that dominates it. The set of the objective values of all Pareto-optimal solutions form the Pareto Front.
@@ 2.3 Benchmarking MORL
Common benchmarks for MORL include game environments, for instance those in Multi-Objective Gymnasium  Felten et al. (2023) such as Deep Sea Treasure (DST)  Vamplew et al. (2011) and Fruit Tree Navigation (FTN)  Yang et al. (2019). These are game environments, e.g. where an agent tries to accomplish multiple objectives in a 2D grid world or MuJoCo task with continuous action spaces  Zhu et al. (2023). However, these benchmarks are still limited to the mechanics of the game environment itself (limited action and observation spaces), and it is difficult to design custom scenarios and objective functions to challenge the RL model. Even though variants of these problems  Cassimon et al. (2021) have also been introduced to increase the size of the observation space and number of objectives, these variants are still non-expandable and normally focus on small action and observation spaces, with limited number of objectives, often just 2-3 objectives  Zhu et al. (2023). GraphAllocBench addresses these gaps by providing a curated benchmark problem set designed using our sandbox environment, CityPlannerEnv.
@@ 4.1 Hypervolume
For each sampled preference $\mathbf{w}$, we obtain a final allocation of resources, which allows us to calculate the final production vector $\mathbf{P}$ for a given allocation. Hence, we can use the final production state to calculate the value of each objective component, which are collectively represented by $\mathbf{J}(\mathbf{P})$. By varying our preferences, we obtain sets of objective vectors $\mathbf{J}$, which form a set of predicted optimal solutions (i.e. the predicted Pareto Front). We can evaluate this solution set is using the hypervolume metric based on the maximization variant of the PyMOO  Blank and Deb (2020) implementation (See Appendix A).
@@ 4.2 Proportion of Non-Dominated Solutions (PNDS)
    We are able to evaluate this metric even if it is computationally infeasible to evaluate the ideal Pareto Front (e.g. high number of production combinations).
@@ 5.1 Main Problem Set
A PCPL agent may find it difficult to deal with sparse rewards, especially when optimizing a continuous preference vector. As example of this would be the $\mathbf{J}_{i}(\mathbf{P})=\text{max}(0,\mathbf{P}_{j}-5)$ floor function, where we receive no reward for Objective $i$ until we produce more than $5$ units for Demand $j$. Instead of deciding whether to a) optimize a particular objective by creating productions until the reward spikes or to b) allocate no resources to the given production, the agent might create insufficient production, such that resources are wasted but no additional reward is received. This results in a lower proportion of non-dominated solutions, as observed for Problem 1c in Figure 2. In Figure 3, we also observe many predicted solutions along the horizontal axis where the reward component for Objective 1 is 0. However, there should only be one Pareto-optimal solution along the horizontal axis that maximizes Objective 0.

### v2

@@ Abstract
Preference-Conditioned Policy Learning (PCPL) in Multi-Objective Reinforcement Learning (MORL) aims to approximate diverse Pareto-optimal solutions by conditioning policies on user-specified preferences over objectives. This enables a single model to flexibly adapt to arbitrary trade-offs at run-time by producing a policy on or near the Pareto front. However, existing benchmarks for PCPL are largely restricted to toy tasks and fixed environments, limiting their realism and scalability. To address this gap, we introduce GraphAllocBench, a flexible benchmark built on a novel graph-based resource allocation sandbox environment inspired by city management, which we call CityPlannerEnv. GraphAllocBench provides a rich suite of problems with diverse objective functions, varying preference conditions, and high-dimensional scalability. We also propose two new evaluation metrics – Proportion of Non-Dominated Solutions (PNDS) and Ordering Score (OS) – that directly capture preference consistency while complementing the widely used hypervolume metric. Through experiments with Multi-Layer Perceptrons (MLPs) and graph-aware models, we show that GraphAllocBench exposes the limitations of existing MORL approaches and paves the way for using graph-based methods such as Graph Neural Networks (GNNs) in complex, high-dimensional combinatorial allocation tasks. Beyond its predefined problem set, GraphAllocBench enables users to flexibly vary objectives, preferences, and allocation rules, establishing it as a versatile and extensible benchmark for advancing PCPL. Code: https://github.com/jzh001/GraphAllocBench
@@ 2.1 Pareto Front
A solution $x^{*}\in X$ is called Pareto optimal if there is no other solution $x\in X$ that dominates it. The set of the objective values of all Pareto-optimal solutions forms the Pareto Front.
@@ 2.3 Benchmarking MORL
Common benchmarks for MORL include game environments, for instance those in Multi-Objective Gymnasium  Felten et al. (2023) such as Deep Sea Treasure (DST)  Vamplew et al. (2011) and Fruit Tree Navigation (FTN)  Yang et al. (2019). These are game environments, e.g. where an agent tries to accomplish multiple objectives in a 2D grid world or MuJoCo task with continuous action spaces  Zhu et al. (2023). However, these benchmarks are still limited to the mechanics of the game environment itself (limited action and observation spaces), and it is difficult to design custom scenarios and objective functions to challenge the RL model. Even though variants of these problems  Cassimon et al. (2021) have also been introduced to increase the size of the observation space and number of objectives, these variants are still non-expandable and normally focus on small action and observation spaces, with a limited number of objectives, often just 2-3 objectives  Zhu et al. (2023). GraphAllocBench addresses these gaps by providing a curated benchmark problem set designed using our sandbox environment, CityPlannerEnv.
@@ 4.1 Hypervolume
For each sampled preference $\mathbf{w}$, we obtain a final allocation of resources, which allows us to calculate the final production vector $\mathbf{P}$ for a given allocation. Hence, we can use the final production state to calculate the value of each objective component, which are collectively represented by $\mathbf{J}(\mathbf{P})$. By varying our preferences, we obtain sets of objective vectors $\mathbf{J}$, which form a set of predicted optimal solutions (i.e. the predicted Pareto Front). We can evaluate this solution set using the hypervolume metric based on the maximization variant of the PyMOO  Blank and Deb (2020) implementation (See Appendix A).
@@ 4.2 Proportion of Non-Dominated Solutions (PNDS)
    We are able to evaluate this metric even if it is computationally infeasible to evaluate the ideal Pareto Front (e.g., a high number of production combinations).
@@ 5.1 Main Problem Set
A PCPL agent may find it difficult to deal with sparse rewards, especially when optimizing a continuous preference vector. An example of this would be the $\mathbf{J}_{i}(\mathbf{P})=\text{max}(0,\mathbf{P}_{j}-5)$ floor function, where we receive no reward for Objective $i$ until we produce more than $5$ units for Demand $j$. Instead of deciding whether to a) optimize a particular objective by creating productions until the reward spikes or to b) allocate no resources to the given production, the agent might create insufficient production, such that resources are wasted but no additional reward is received. This results in a lower proportion of non-dominated solutions, as observed for Problem 1c in Figure 2. In Figure 3, we also observe many predicted solutions along the horizontal axis where the reward component for Objective 1 is 0. However, there should only be one Pareto-optimal solution along the horizontal axis that maximizes Objective 0.